OpenStudy (anonymous):

Can anyone help to simplfy. 1 over 3x+9 minus 2 over 2x3=2

OpenStudy (anonymous):

draw it !!

OpenStudy (dumbcow):

what do you mean by " 2x3" ?

OpenStudy (dumbcow):

is it \[\frac{1}{3x+9} - \frac{2}{2*3} = 2\]

OpenStudy (dumbcow):

this looks better....2/6 = 1/3 factor out a 3 from 3x+9 \[\rightarrow \frac{1}{3(x+3)}-\frac{1}{3} = 2\] Factor out a 1/3 \[\rightarrow \frac{1}{3}(\frac{1}{x+3} -1) = 2\] \[\rightarrow \frac{1}{x+3} = 7\] \[1 = 7x +21\] \[x = -\frac{20}{7}\]